HETAV PATEL

Grade 11,

What is fajan's rule?/////////////////////////////////////////////////////

What is fajan's rule?/////////////////////////////////////////////////////

Grade:11

2 Answers

Shivangi Khatter
5 years ago
dear student
Fajans' rules, formulated byKazimierz Fajansin 1923,[1][2][3]are used to predict whether achemical bondwill becovalentorionic, and depend on the charge on thecationand the relative sizes of the cation andanion

also,Thussodium chloride(with a low positive charge (+1), a fairly large cation (~1Å) and relatively small anion (0.2Å) is ionic; butaluminium iodide(AlI3) (with a high positive charge (+3) and a large anion) is covalent.

Polarization will be increased by:

high charge and small size of the cation

Ionic potential Å Z+/r+ (= polarizing power)

High charge and large size of the anion

The polarizability of an anion is related to the deformability of its electron cloud (i.e. its "softness")

An incomplete valence shell electron configuration

Noble gas configuration of the cation produces better shielding and less polarizing power

e.g. Hg2+(r+ = 102 pm) is more polarizing than Ca2+(r+ = 100 pm)
Khimraj
5 years ago

They are a method for predicting ionic vs. covalent that predates electronegativity (by three decades) and make use of ionic and atomic radius data that was becoming available through x-ray crystallography.

To use Fajans' Rules, assume your binary compound is ionic and identify the potential cation and anion.

By Fajans' Rules, compounds are more likely to be ionic if: there is a small positive charge on the cation, the cation is large, and the anion is small. For example, NaCl is correctly predicted to be ionic since NaX+  is a larger ion with a low charge and ClX- is a smaller anion.

Compounds are more likely to be covalent if: there would be a large positive charge on the cation, the cation would be small, and the anion would be large. For example, AlI3is correctly predicted to be covalent since it would have a small cation with a high charge and a large anion